OpenStudy (zarakkhalil):

Two competing cell phone companies have recently restructured their payment plans. Mobilama requires a $20 activation fee and charges $25 per month for their basic plan. Ringit Wireless requires no activation fee and charges $35 per month for their basic plan. When will the two plans be equal in cost?

OpenStudy (sooobored):

we can write these two payment plans in the form of a function the total cost of mobilama m(x) and the total cost of ringit r(x) m(x) = 20+25x -- start with a 20 activation fee, and every month "x" we add an additional 25 r(x) = 0 + 35x -- no activation fee, and every month "x" we add an additional 35 to the bill

OpenStudy (sooobored):

we want to determine at what time/ how many month will the total costs of these 2 payment plans be equal to each other so when m(x)=r(x) thus 20+25x=35x and then solve for x which is hte total amount of months required before both payment plans have the same total cost